Who When
M. Bonaglia 20140506 - 20140520
L. Busoni 20140506 - 20140520
T. Mazzoni 20140506 - 20140520
M. Kulas 20140508 - 20140520
J. Borelli 20140508 - 20140520
J. Ziegleder 201405?? - 20140520
G. Orban de Xivry 20140506 - 20140520
W. Raab 201405?? - 20140520
D. Peter 20140501 - 20140507
M. Lefebvre 20140501 - 20140508
G. Rahmer 201405?? - 201405??
W. Gaessler 20140501 - 20140520
S. Rabien 201405?? - 20140520
S. Esposito 201405?? - 20140520


2014_03_ActionItems_v3.pdf: Action Items from Mar 13 run


Activity summary

When What Log Notes
20140501 Thu Cal unit activity. DayTime20140501  
20140502 Fri Cal unit activity. DayTime20140502  
20140503 Sat Cal unit activity. DayTime20140503  
20140504 Sun Cal unit activity. DayTime20140504  
20140505 Mon Cal unit activity. DayTime20140505  
20140506 Tue Cal unit activity. DayTime20140506  
20140507 Wed Cal unit activity. IM software developement. DayTime20140507 ComaCorrection
20140508 Thu Installed filter/regulator on PnCCD cooling line, cable cover and roll bar. Pockels cell test with portable laser. No cal unit available. DayTime20140508  
20140509 Fri AdSec Pie-Shape during morning. Cal unit deployed at 11:50. Im recording and test with 10, 50, 100 modes. PockelsCellTest on Yellow pupil. DayTime20140509 CalUnitTricks
20140510 Sat Calibration of yellow and blue pupils up to 150 modes. DayTime20140510    
20140511 Sun Calibration of red pupil uo to 50 modes. FLAO CL on on-axis source, LUCI images. ARGOS CL on yellow and blue source, LUCI images. Laser power measured and OK. Activity on SX lasers in the lab. DayTime20140511  
20140512 Mon AdSec Pie-Shape during night, recovers about 9 a.m. We try to close the TT loop with FLAO and HO with ARGOS blue pupil. Fail. AdSec Pie-Shape again at 12 a.m. recovers at 5 p.m. We try to close the FLAO TT loop and ARGOS HO on blue pupil. Success! PockelsCellTest on Yellow during evening. DayTime20140512  
20140513 Tue AdSec Pie-Shape. M.Wagner try to disconnect the accelerometer board. PockelsCellTest on Red and Blue. Check of the BCU serial tip-tilt and FLAO communications. DayTime20140513  
20140514 Wed Beginning of NightTime operations. AdSec still pieshaped. Activity on the LGS acquisition (IMs: PerPatCam, LM1PatCam, PerQC, PupQC). LanLookupTable. NightTime20140514  
20140515 Thu PockelsCellTest on Blue and SX PC #5. AdSec recovers. IM and REC on yellow (150modes). ARGOS + PYR loop. SlopesNull acquisition. LAN focus scan. ARGOS CL. NightTime20140515  
20140516 Fri Cal unit activity. Cloudy night. NightTime20140516  
20140517 Sat Software development. DayTime: OL and CL ARGOS test (up to 10modes). NightTime: test of LAN optical quality test looking at NGS, LanLookupTable. DayTime20140517 NightTime20140517  
20140518 Sun DayTime: new calibrations on yellow. NightTime: clouds during first part of night, LAN optical quality test, CL ARGOS test in late night (50modes). DayTime20140518 NightTime20140518 Report_On-sky_Vignetting.docx
20140519 Mon PockelsCells dismount DayTime20141119  
Topic revision: r18 - 24 Sep 2015, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback